# Multimodal foundation world models for generalist embodied agents

### v1

@@ 4.1 Offline RL
Table 1: Behavior extraction. Offline RL from language prompts on tasks that are included in the agent’s training dataset. Scores are episodic rewards averaged over 5 seeds (± standard error) rescaled using min-max scaling with $(\min=\textrm{random policy},\max=\textrm{expert policy})$.
Image-language reward Video-language reward Ours
IQL TD3+BC TD3 IQL TD3+BC TD3 GenRL
walker stand 0.68 ± 0.03 0.97 ± 0.01 0.92 ± 0.06 0.72 ± 0.05 0.59 ± 0.05 1.0 ± 0.0 1.01 ± 0.0
walker run 0.24 ± 0.03 0.27 ± 0.02 0.11 ± 0.03 0.26 ± 0.02 0.25 ± 0.02 0.35 ± 0.01 0.76 ± 0.01
walker walk 0.37 ± 0.04 0.25 ± 0.03 0.15 ± 0.0 0.43 ± 0.04 0.39 ± 0.01 0.86 ± 0.04 0.95 ± 0.03
cheetah run 0.37 ± 0.07 -0.01 ± 0.0 -0.01 ± 0.0 0.15 ± 0.03 -0.01 ± 0.0 0.34 ± 0.01 0.74 ± 0.01
quadruped stand 0.32 ± 0.06 0.4 ± 0.06 0.62 ± 0.04 0.39 ± 0.04 0.41 ± 0.13 0.73 ± 0.06 0.98 ± 0.0
quadruped run 0.3 ± 0.03 0.35 ± 0.01 0.25 ± 0.02 0.36 ± 0.04 0.35 ± 0.04 0.23 ± 0.02 0.61 ± 0.04
quadruped walk 0.16 ± 0.02 0.18 ± 0.02 0.37 ± 0.02 0.19 ± 0.04 0.16 ± 0.04 0.37 ± 0.02 0.73 ± 0.06
stickman run 0.2 ± 0.02 0.25 ± 0.02 0.03 ± 0.0 0.24 ± 0.02 0.18 ± 0.03 0.21 ± 0.0 0.35 ± 0.02
stickman walk 0.43 ± 0.07 0.52 ± 0.05 0.18 ± 0.01 0.47 ± 0.02 0.45 ± 0.08 0.42 ± 0.03 0.78 ± 0.02
stickman stand 0.52 ± 0.05 0.68 ± 0.04 0.56 ± 0.04 0.56 ± 0.06 0.47 ± 0.02 0.61 ± 0.03 0.49 ± 0.04
kitchen microwave 0.14 ± 0.13 0.0 ± 0.0 0.0 ± 0.0 0.02 ± 0.02 0.0 ± 0.0 0.46 ± 0.18 0.67 ± 0.27
kitchen light 0.18 ± 0.12 1.0 ± 0.0 0.66 ± 0.07 0.04 ± 0.02 0.0 ± 0.0 0.0 ± 0.0 0.43 ± 0.19
kitchen burner 0.14 ± 0.07 0.43 ± 0.14 0.02 ± 0.02 0.02 ± 0.02 0.0 ± 0.0 0.28 ± 0.08 0.67 ± 0.14
kitchen slide 0.16 ± 0.1 0.0 ± 0.0 0.0 ± 0.0 0.06 ± 0.04 0.0 ± 0.0 1.0 ± 0.0 1.0 ± 0.0
overall 0.30 ± 0.04 0.38 ± 0.02 0.28 ± 0.02 0.28 ± 0.02 0.23 ± 0.02 0.49 ± 0.04 0.73 ± 0.05
@@
Behavior extraction. We want to verify whether the methods can retrieve the tasks behaviors that are certainly present in the dataset. We present results in Table 1, with episodic rewards rescaled so that 0 represents the performance of a random agent, while 1 represents the performance of an expert agent.
@@
GenRL excels in overall performance across all domains, particularly in dynamic tasks like walking and running in the quadruped and cheetah domains. In contrast, in static tasks such as ’stickman stand’ and kitchen tasks, other methods occasionally outperform GenRL. This can be explained by the fact that the target sequences that GenRL infers from the prompt are often slightly in motion, even in static cases. To address this, we could set the target sequence length to 1 for static prompts, but we opted to maintain the method’s simplicity and generality, acknowledging this as a minor limitation.
@@
As expected, video-language rewards tend to perform better than image-language rewards for dynamic tasks. For video-based rewards, the less conservative approach, TD3, performs better than all other baselines in most tasks, similarly to what is shown in [58]. We instead observe an opposite trend for image-language rewards, where more conservative approaches, such as IQL and TD3+BC, tend to perform better. We believe this is because when the ‘task target’ is static, imitating segments of trajectories from the dataset proves beneficial.
@@
Figure 3: Multi-task generalization. Offline RL from language prompts on tasks that are not deliberately included in the training dataset. Performance averaged over 5 seeds and standard error was reported with black lines. Detailed results per task in Appendix E.
@@
Multi-task generalization. To assess multi-task generalization, we defined a set of tasks not included in the training data. Although we don’t anticipate agents matching the performance of expert models, higher scores in this benchmark help gauge the generalization abilities of different methods. We averaged the performance across various tasks for each domain and summarized the findings in Figure 3, with detailed task results in Appendix E.
@@
Overall, we observe a similar trend as for the behavior extraction results. GenRL significantly outperforms the other approaches, especially in the quadruped and cheetah domains, where the performance is close to the specialized agents’ performance. Both for image-language (-I in the Figure) and video-language (-V in the Figure) more conservative approaches, such as IQL and TD3+BC tend to perform worse. This could be associated with the fact that imitating segments of trajectories is less likely to lead to high-rewarding trajectories, as the tasks are not present in the training data.
4.2 Data-free RL
@@
Figure 4: Data-free RL. Generalization to new tasks, learning behaviors in imagination without relying on any data for initializing the sequences to learn in imagination. Performance is averaged over 5 seeds and standard error is reported with black lines. Detailed results per task in Appendix E.
@@
In the previous section, we evaluated several approaches for designing reward using foundation VLMs. Clearly, model-free RL approaches require continuous access to a dataset, to train the actor-critic and generalize across new tasks. Model-based RL can learn the actor-critic in imagination. However, in previous work [25, 23], imagining sequences for learning behaviors first requires processing actual data sequences. The data is used to initialize the dynamics model, and obtain latent states that represent the starting states to rollout the policy in imagination. Furthermore, in order to learn new tasks, reward-labelled data is necessary to learn a reward model, which provides rewards to the agent during the task learning process.
@@
Foundation models [41] are generally trained on enormous datasets in order to generalize to new tasks. The datasets used for the model pretraining are not necessary for the downstream applications, and sometimes these datasets are not even publicly available [40, 18]. In this section, we aim to establish a new paradigm for foundation models in RL, which follows the same principle of foundation models for vision and language. We call this paradigm data-free RL and we define it as the ability to generalize to new tasks, after pre-training, using no additional data.
@@
GenRL enables data-free RL thanks to two main reasons: the agent learns a task-agnostic MFWM on a large varied dataset during pre-training, and the MFWM enables the possibility of specifying tasks directly in latent space, without requiring any data. Thus, in order to learn behaviors in imagination, the agent can: (i) sample random latent states in the world model’s representation, (ii) rollout sequences in imagination, following the policy, and (iii) compute rewards, using the targets obtained by processing the given prompts with the connector-aligner networks.
Performance of data-free RL. Our results, detailed in Figure 4, compare data-free RL to traditional offline RL with GenRL, as discussed in Section 4.1. We ablate the choice of using random samples from the connector-aligner to improve randomly sampled initial states. While data-free RL generally shows a slight decrease in overall performance, the differences are minimal across most domains, and it even outperforms in the kitchen domain. The use of states from the connector model enhances average scores and reduces variance, especially noticeable in the cheetah domain.
@@
By employing data-free learning, after pre-training, agents can master new tasks without data, often converging within only 30 minutes of GPU training. As we scale up foundation models for behavior learning, the ability to learn data-free will become crucial. Although very large datasets will be employed to train new foundation models, GenRL adapts well without direct access to original data, offering flexibility where data may be proprietary, licensed or unavailable.
4.3 Training data distribution
@@
Figure 5: Training data distribution. Analysing the impact of the training data distribution on the generalization performance of GenRL. Performance is obtained by training behaviors in data-free mode, after training the MFWM on different subsets of the training dataset. Performance averaged over 3 seeds (black lines indicate standard error). Full results in Appendix E.
@@
As demonstrated in Sections 4.1 and 4.2, after training on a large dataset, a GenRL agent can adapt to multiple new tasks without additional data. The nature of the training data, detailed in Appendix C, combines exploration and task-specific data. To identify critical data types for GenRL, we trained different MFWMs on various dataset subsets. Then, we employ data-free RL to train task behaviors, with analyses over subsets of the walker dataset provided in Figure 5, where ‘all’ reports the data-free performance when training on the full dataset.
@@
The results confirm that a diverse data distribution is crucial for task success, with the best performance achieved by using the complete dataset, followed by the varied exploration data. Task-specific data effectiveness depends on task complexity, for instance, ’run’ data proves more generalizable than ’walk’ or ’stand’ data across tasks. Crucially, ’stand’ data, which shows minimal variation, limits learning for a general agent but can still manage simpler tasks like ’lying down’ and ’sitting on knees’ as detailed in Appendix E.
@@
5 Additional Analysis
A framework for behavior generation. A common challenge with using LLMs and VLMs involves the need for prompt tuning to achieve specific tasks. As GenRL relies on a foundation VLM, similar to previous approaches [4, 45] it is not immune from this issue. However, GenRL uniquely allows for the visualization of targets obtained from specific prompts. By decoding the latent targets, using the MFWM decoder, we can visualize the interpreted prompt before training the corresponding behavior. This enables a much more explainable framework, which allows fast iteration for prompt tuning, compared to previous (model-free) approaches which often require training the agent to identify which behaviors are rewarded given a certain prompt.
Video grounding. In Section 4, we have focussed our evaluation on language-driven RL, using text prompts to specify tasks. While language strongly simplifies specification of the task, in some cases providing visual examples of the task might be easier. Similarly as for language prompts, GenRL allows translating vision prompts (short videos) into behaviors. As this is a less common evaluation setting, we limit our assessment to qualitative results.
In Figure 6(a), we present a set of video grounding examples, obtained by inferring the latent targets corresponding to the vision prompts (right image) and then using the decoder model to decode images (left image). We observe that the agent is able to translate short human action videos [^2] Short videos are generated using meta.ai. into the same actions but for the Stickman embodiment. By applying this approach, it would be possible to learn behaviors from a single video.
Scaling to complex observations. Generalist embodied agents should be able to scale to open-ended learning settings. Using GenRL, we explored this by training an agent in the Minecraft environment using a small dataset collected by a DreamerV3 agent [25]. The primary challenge we found was the model’s difficulty in reconstructing complex observations in this open-ended environment.
Reconstructing complex observations is a common issue with world models [10]. To overcome this limitation, while keeping the method unaltered, we attempted to scale up the number of parameters of MFWM. Qualitative reconstruction results are presented in Figure 6(b). We observe that the agent is able to identify different biomes from language, even with the smaller size of the model. However, the reconstructions are significantly blurrier compared to the other environments we analyzed (e.g. Figure 6(a)). When using a larger model, the reconstructions gain some details but the results still highlight the difficulty of the model in providing accurate targets from prompts.
While this might not be an issue for simple high-level tasks, e.g. ‘navigate to a beach’, unclear targets might make it difficult to perform more precise actions, e.g. ‘attack a zombie’. Future research should aim to address this issue, for instance, by improving our simple GRU-based architecture, leveraging transformers or diffusion models to improve the quality of the representation [29, 3].
(a) Video grounding examples
(b) Decoded language prompts in Minecraft
@@ 6 Discussion
We introduced GenRL, a world-model based approach for grounding vision-language prompts into embodied domains and learning the corresponding behaviors in imagination. The multimodal foundation world models of GenRL can be trained using unimodal data, overcoming the lack of multimodal data in embodied RL domains. The data-free RL property of GenRL lays the groundwork for foundation models in RL that can generalize to new tasks without additional data.
Limitations. Despite its strengths, GenRL presents some limitations, largely due to inherent weaknesses in its components. From the VLMs, GenRL inherits the issue related to the multimodality gap [33, 61] and the reliance on prompt tuning. We proposed a connection-alignment mechanism to mitigate the former. For the latter, we presented an explainable framework, which facilitates prompt tuning by allowing decoding of the latent targets corresponding to the prompts. From the world model, GenRL inherits a dependency on reconstructions, which offers advantages such as explainability but also drawbacks, such as failure modes with complex observations.

### v2

@@ 4.1 Offline RL
Table 1: Language-to-action in-distribution. Offline RL from language prompts on tasks that are included in the agent’s training dataset. Scores are episodic rewards averaged over 10 seeds (± standard error) rescaled using min-max scaling with $(\min=\textrm{random policy},\max=\textrm{expert policy})$.
Image-language VLM Video-language VLM
IQL TD3+BC TD3 WM-CLIP IQL TD3+BC TD3 WM-CLIP GenRL
walker stand 0.67 ± 0.03 0.92 ± 0.02 0.93 ± 0.03 1.01 ± 0.0 0.66 ± 0.05 0.64 ± 0.03 1.01 ± 0.0 0.94 ± 0.01 1.02 ± 0.0
walker run 0.24 ± 0.03 0.27 ± 0.01 0.09 ± 0.02 0.05 ± 0.02 0.29 ± 0.02 0.24 ± 0.02 0.35 ± 0.01 0.7 ± 0.01 0.77 ± 0.02
walker walk 0.41 ± 0.05 0.34 ± 0.05 0.14 ± 0.0 0.21 ± 0.01 0.4 ± 0.03 0.44 ± 0.03 0.88 ± 0.02 0.91 ± 0.02 1.01 ± 0.0
cheetah run 0.41 ± 0.05 0.0 ± 0.01 -0.01 ± 0.0 -0.0 ± 0.0 0.15 ± 0.02 -0.01 ± 0.0 0.37 ± 0.01 0.56 ± 0.03 0.74 ± 0.01
quadruped stand 0.56 ± 0.02 0.64 ± 0.04 0.65 ± 0.04 0.97 ± 0.0 0.52 ± 0.06 0.43 ± 0.05 0.61 ± 0.05 0.97 ± 0.0 0.97 ± 0.0
quadruped run 0.3 ± 0.03 0.28 ± 0.02 0.24 ± 0.02 0.27 ± 0.0 0.38 ± 0.03 0.25 ± 0.02 0.26 ± 0.01 0.61 ± 0.02 0.86 ± 0.02
quadruped walk 0.26 ± 0.02 0.31 ± 0.02 0.28 ± 0.01 0.47 ± 0.02 0.32 ± 0.02 0.28 ± 0.04 0.28 ± 0.02 0.92 ± 0.01 0.93 ± 0.01
stickman stand 0.45 ± 0.06 0.58 ± 0.04 0.06 ± 0.04 0.71 ± 0.02 0.43 ± 0.04 0.45 ± 0.05 0.08 ± 0.02 0.32 ± 0.01 0.7 ± 0.02
stickman walk 0.4 ± 0.04 0.48 ± 0.04 0.18 ± 0.01 0.23 ± 0.01 0.51 ± 0.02 0.46 ± 0.03 0.41 ± 0.02 0.65 ± 0.05 0.83 ± 0.01
stickman run 0.2 ± 0.01 0.22 ± 0.02 0.03 ± 0.0 0.19 ± 0.01 0.23 ± 0.02 0.19 ± 0.02 0.21 ± 0.0 0.35 ± 0.01 0.35 ± 0.01
kitchen microwave 0.06 ± 0.04 0.22 ± 0.11 0.0 ± 0.0 0.0 ± 0.0 0.01 ± 0.01 0.0 ± 0.0 0.11 ± 0.08 0.9 ± 0.09 0.97 ± 0.02
kitchen light 0.14 ± 0.04 0.11 ± 0.11 0.59 ± 0.16 0.1 ± 0.09 0.02 ± 0.01 0.0 ± 0.0 0.18 ± 0.11 0.26 ± 0.13 0.46 ± 0.09
kitchen burner 0.21 ± 0.05 0.18 ± 0.05 0.09 ± 0.05 0.03 ± 0.03 0.05 ± 0.02 0.02 ± 0.01 0.31 ± 0.1 0.78 ± 0.06 0.62 ± 0.07
kitchen slide 0.02 ± 0.01 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.04 ± 0.02 0.02 ± 0.02 0.7 ± 0.14 0.88 ± 0.04 1.0 ± 0.0
overall 0.31 ± 0.03 0.33 ± 0.04 0.23 ± 0.04 0.30 ± 0.02 0.29 ± 0.02 0.24 ± 0.02 0.41 ± 0.05 0.70 ± 0.04 0.80 ± 0.02
@@
Language-to-action in-distribution. We want to verify whether the methods can retrieve the task behaviors that are certainly present in the training data, when specifying the task only through language. We present results in Table 1, with episodic rewards rescaled so that 0 represents the performance of a random agent, while 1 represents the performance of an expert agent.
@@
GenRL excels in overall performance across all domains and tasks, outperforming other methods particularly in dynamic tasks like walking and running in the quadruped and cheetah domains. However, in some static tasks of the kitchen domain, other methods occasionally outperform GenRL. This can be explained by the fact that the target sequences that GenRL infers from the prompt are often slightly in motion, even in static cases. To address this, we could set the target sequence length to 1 for static prompts, but we opted to maintain the method’s simplicity and generality, acknowledging this as a minor limitation.
@@
As expected, video-language rewards tend to perform better than image-language rewards for dynamic tasks. The less conservative approach, TD3, performs better than the other model-free baselines in most tasks, similarly to what is shown in Yarats2022ExORL . The model-based baseline’s performance, WM-CLIP-V, is the closest to GenRL’s.
@@
Figure 4: Language-to-action generalization. Offline RL from language prompts on tasks that are not deliberately included in the training dataset. Performance averaged over 10 seeds and standard error was reported with black lines. Detailed results per task in Appendix K.
@@
Language-to-action generalization. To assess multi-task generalization, we defined a set of tasks not included in the training data. Although we don’t anticipate agents matching the performance of expert models, higher scores in this benchmark help gauge the generalization abilities of different methods. We averaged the performance across various tasks for each domain and summarized the findings in Figure 4, with detailed task results in Appendix K.
@@
Overall, we observe a similar trend as for the in-distribution results. GenRL significantly outperforms all model-free approaches, especially in the quadruped and cheetah domains, where the performance is close to the specialized agents’ performance. Both for image-language (-I in the Figure) and video-language (-V in the Figure) more conservative approaches, such as IQL and TD3+BC tend to perform worse. This could be associated with the fact that imitating segments of trajectories is less likely to lead to high-rewarding trajectories, as the tasks are not present in the training data.
Figure 5: Video-to-action. GenRL allows grounding video prompts into the target environment’s dynamics. It allows visualization of the model’s interpretation of the prompts, using the decoder (top row), and it allows turning prompts into behaviors, leading to generally higher performance than other approaches. 10 seeds. Additional visualizations on the project website.
Video-to-action. While language strongly simplifies the specification of a task, in some cases providing visual examples of the task might be easier. Similarly as for language prompts, GenRL allows grounding visual prompts (short videos) into the embodied domain’s dynamics and then learning of the corresponding behaviors.
In Figure 5, we provide behavior learning results from video prompts. The tasks included are of static and dynamic nature and span across 4 different domains. Visualizations of the videos used as prompts are available on the project website, where we also present a set of “grounded videos" generated by the model using the prompts (see snapshots at the top of Fig. 5). These can be obtained by inferring the latent targets corresponding to the vision prompts (left images, in the Figure) and then using the decoder model to decode reconstructed images (right images, in the Figure).
The results show a similar trend to the language prompts experiments and the performance when using video prompts is aligned to the language-to-action performance, for the same tasks. In general, we found it interesting that the VLM allows us to generalize to very different visual styles (drawings, realistic, AI-generated), very different camera viewpoints (quadruped, microwave), and different morphologies (cheetah tasks).
Summary. The experiments presented allow us to establish more clearly the main ingredients that contribute to the stronger performance of GenRL: (i) the video-language model helps in dynamic tasks, (ii) model-based algorithms lead to higher performance, (iii) the connection-alignment system presented generally outperforms the “reversed" way of connecting the two representations.
4.2 Data-free policy learning
@@
Figure 6: By removing data dependencies on actions (on-policy learning in the model’s imagination), rewards (computed using only the prompt and latent states, using Eq. 3), and observations (by sampling latent states within the model), GenRL agents can be adapted for new tasks in a data-free fashion. Performance is averaged over 10 seeds and standard error is reported with black lines. Detailed results per task in Appendix K.
@@
In the previous section, we evaluated several approaches for designing reward using foundation VLMs. Clearly, model-free RL approaches require continuous access to a dataset, to train the actor-critic and generalize across new tasks. Model-based RL can learn the actor-critic in imagination. However, in previous work Hafner2023DreamerV3 ; Hafner2020DreamerV1 , imagining sequences for learning behaviors first requires processing actual data sequences. The data is used to initialize the dynamics model, and obtain latent states that represent the starting states to rollout the policy in imagination. Furthermore, in order to learn new tasks, reward-labelled data is necessary to learn a reward model, which provides rewards to the agent during the task learning process.
@@
Foundation models Bommasani2022FoundationModels are generally trained on enormous datasets in order to generalize to new tasks. The datasets used for the model pretraining are not necessary for the downstream applications, and sometimes these datasets are not even publicly available OpenAI2023GPT4 ; Google2024Gemini . In this section, we aim to establish a new paradigm for foundation models in RL, which follows the same principle of foundation models for vision and language. We call this paradigm data-free policy learning and we define it as the ability to generalize to new tasks, after pre-training, by learning a policy completely in imagination, with no access to data (not even to the pre-training dataset).
@@
GenRL enables data-free policy learning thanks to two main reasons: the agent learns a task-agnostic MFWM on a large varied dataset during pre-training, and the MFWM enables the possibility of specifying tasks directly in latent space, without requiring any data. Thus, in order to learn behaviors in imagination, the agent can: (i) sample random latent states in the world model’s representation, (ii) rollout sequences in imagination, following the policy’s actions, and (iii) compute rewards, using the targets obtained by processing the given prompts with the connector-aligner networks.
In Figure 6, we provide a diagram that further clarifies the differences between training GenRL in an offline and in a data-free policy learning fashion. Then, we present results that compare data-free policy learning with offline RL baselines, as discussed in Section 4.1. While data-free policy learning shows a slight decrease in overall performance, its performance remains close to the original GenRL’s performance and still outperforms other approaches. In Appendix K, we further show that the difference in performance is minimal across most domains, and data-free policy learning even performs better in the kitchen domain.
@@
By employing data-free learning, after pre-training, agents can master new tasks without data. By requiring no CPU-GPU memory transfers of the data, data-free policy learning also reduces the training time of the policy, often allowing convergence within only 30 minutes of training. As we scale up foundation models for behavior learning, the ability to learn data-free will become crucial. Although very large datasets will be employed to train future foundation models, GenRL adapts well without direct access to original data, offering flexibility where data may be proprietary, licensed or unavailable.
4.3 Analysis of the training data distribution
@@
Figure 7: Training data distribution. Analysing the impact of the training data distribution on the generalization performance of GenRL. Performance is obtained by training behaviors in data-free mode, after training the MFWM on different subsets of the training dataset. Performance averaged over 10 seeds (black lines indicate standard error). Full results in Appendix K.
@@
As demonstrated in Sections 4.1 and 4.2, after training on a large dataset, a GenRL agent can adapt to multiple new tasks without additional data. The nature of the training data, detailed in Appendix C, combines exploration and task-specific data. Thus, we ask ourselves what subsets of the data are the most important ones for GenRL’s training.
To identify critical data types for GenRL, we trained different MFWMs on various dataset subsets. Then, we employ data-free behavior learning to train task behaviors for all tasks. We present an analysis over subsets of the walker dataset in Figure 7.
@@
The results confirm that a diverse data distribution is crucial for task success, with the best performance achieved by using the complete dataset, followed by the varied exploration data. Task-specific data effectiveness depends on task complexity, for instance, ’run data’ proves more useful and generalizable than ’walk data’ or ’stand data’ across tasks. Crucially, ’stand data’, which shows minimal variation, limits learning for a general agent but can still manage simpler tasks like ’lying down’ and ’sitting on knees’ as detailed in Appendix K.
@@ 5 Discussion
We introduced GenRL, a world-model based approach for grounding vision-language prompts into embodied domains and learning the corresponding behaviors in imagination. The multimodal-foundation world models of GenRL can be trained using unimodal data, overcoming the lack of multimodal data in embodied RL domains. The data-free behavior learning capacity of GenRL lays the groundwork for foundation models in RL that can generalize to new tasks without any data.
A framework for behavior generation. A common challenge with using LLMs and VLMs involves the need for prompt tuning to achieve specific tasks. As GenRL relies on a foundation VLM, similar to previous approaches Baumli2023VLMRew ; Rocamonde2023VLMRM it is not immune from this issue. However, GenRL uniquely allows for the visualization of targets obtained from specific prompts. By decoding the latent targets, using the MFWM decoder, we can visualize the interpreted prompt before training the corresponding behavior. This enables a much more explainable framework, which allows fast iteration for prompt tuning, compared to previous (model-free) approaches which often require training the agent to identify which behaviors are rewarded given a certain prompt.
Limitations. Despite its strengths, GenRL presents some limitations, largely due to inherent weaknesses in its components. From the VLMs, GenRL inherits the issue related to the multimodality gap Liang2022MMGap ; Zhang2024C3 and the reliance on prompt tuning. We proposed a connection-alignment mechanism to mitigate the former. For the latter, we presented an explainable framework, which facilitates prompt tuning by allowing decoding of the latent targets corresponding to the prompts. From the world model, GenRL inherits a dependency on reconstructions, which offers advantages such as explainability but also drawbacks, such as failure modes with complex observations. We further investigate this limitation in Appendix I and present other potential limitations in Appendix J.